OpenStudy (anonymous):

A playground is surrounded by a fence 2.0 m high. A boy who plays in the playground kicks a ball vertically upward from a height of 0.2 m with a speed of 10 m/s. How much time does the ball spend above the height of the fence? (

OpenStudy (anonymous):

a=-g v=-gt+v0 when t=0, v=10 v=-gt+10 x=-g/2 t^2+10t+x0 when t=0, x=0.2 x=-g/2 t^2+10t+0.2 when x=2 2=-g/2 t^2+10t+0.2 0=-g/2 t^2+10t-1.8 0=t^2-20/g t+ 3.6/g 0=t^2-2.04t+0.37 Using the quadratic rule, t=0.20 and t=1.84 So the ball stays in the air for 1.64 seconds.

OpenStudy (amistre64):

i can never remember the physics equation and always resort to the parabola :/

OpenStudy (amistre64):

that is the parabola :) yay!!